OpenStudy (anonymous):

Find the slope of the line passing through 3 -2 and 6 -4?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

with your eyes: from 3 to 6 in the x is over three from -2 to -4 in the y is down two. over three down two means over one down 2/3 so -2/3

OpenStudy (anonymous):

you can always compute \[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

which in this case would be \[\frac{-4-(-2)}{6-3}=\frac{-4+2}{3}=-\frac{2}{3}\]